# Resource-Efficient Beam Prediction in mmWave Communications with Multimodal Realistic Simulation Framework

### v1

@@ VI Conclusion
In this paper, we have developed a CRKD framework for efficient mmWave beam prediction, in which a multimodal teacher model (LiDAR, radar, GPS, and RGB) transfers relational knowledge to a radar-only student. To achieve realistic evaluations, we have integrated CARLA-based sensor data generation with MATLAB-based mmWave channel modeling, creating a comprehensive simulation environment. Experimental results in 2-Lane and 3-Lane road scenarios confirmed that a compact student model with radar can only approach the performance of a transformer-based teacher rich in sensors. This indicates that real-world applications remain viable even with fewer sensor modalities and a smaller network structure. In future work, we plan to incorporate domain adaptation techniques into the CRKD process to further enhance the generalization of the student model under changing conditions. We will also integrate additional sensors and more complex mobility patterns into our simulation framework, with the aim of maintaining high beam prediction accuracy in resource-constrained deployment scenarios and across diverse real-world environments.

### v2

@@ VI Conclusion
In this paper, we have developed a novel CRKD framework for efficient mmWave beam prediction, in which a multimodal teacher model (LiDAR, radar, GPS, and RGB) transfers relational knowledge to a radar-only student. To achieve realistic evaluations, we have integrated CARLA-based sensor data generation with MATLAB-based mmWave channel modeling, creating a comprehensive simulation environment. Experimental results in 2-Lane and 3-Lane road scenarios confirmed that a compact student model with radar can only approach the performance of a transformer-based teacher rich in sensors. This indicates that real-world applications remain viable even with fewer sensor modalities and a smaller network structure. In future work, we plan to incorporate domain adaptation techniques into the CRKD process to further enhance the generalization of the student model under changing conditions. We will also integrate additional sensors and more complex mobility patterns into our simulation framework, with the aim of maintaining high beam prediction accuracy in resource-constrained deployment scenarios and across diverse real-world environments.